OpenStudy (anonymous):

A company offers a starting yearly salary of 37,000 with rises of 2,500 per year. Find the total salary over a ten-tear period

OpenStudy (anonymous):

I think it's $395,00

OpenStudy (anonymous):

$395,000

OpenStudy (anonymous):

I got 59,500

OpenStudy (anonymous):

I got...

OpenStudy (anonymous):

Yeah I think it's that and not 62,000 if someone thinks that

OpenStudy (anonymous):

ok I did something wrong gimmie a minute

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Brb

OpenStudy (anonymous):

ok I got 62,000

OpenStudy (jim766):

an = a1+(n-1)d a10 = 37000 + (10-1)2500 a10 = 59,500 , that is how much he would make in 10th year To find the sum of a arith seq sn = n(a1+an)/2 sn = 10((37000+59500)/2 ) s10 = 482,500

OpenStudy (anonymous):

that's what i got @Jim766

OpenStudy (anonymous):

Yeah I don't know if it is 59,500, 62,000 or another answer

OpenStudy (anonymous):

was 59,500

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

if you multiply 2500 x the ten years then add 37,000

OpenStudy (anonymous):

No wrong formula

OpenStudy (anonymous):

for me, that is

OpenStudy (anonymous):

I thought of a wrong formula, not anyone else

OpenStudy (jim766):

salary in10th year 59,500 but that is not the salary of all 10 years put together.

OpenStudy (anonymous):

Yeah,

OpenStudy (anonymous):

@satellite73 can you clear this up?

OpenStudy (anonymous):

do you have to do that due it asking for the total amount of a 10 year salary?

OpenStudy (anonymous):

Be right back

OpenStudy (jim766):

I think that when it says the salary over the whole period, they want all the salaries added together for each year. Key work is "over" the 10 years

OpenStudy (anonymous):

To me that doesn't make sense because that isn't what they are really making is it?

OpenStudy (anonymous):

@Directrix insight?

OpenStudy (anonymous):

Yeah, that is correct

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

"Total" salary

OpenStudy (anonymous):

Meaning all added up

OpenStudy (anonymous):

can you explain why it's all added up. I am curious not trying to be a pain. LOL

OpenStudy (anonymous):

I think they are asking what is the amount of salary change over 10 years, then add that change to the first year, but don't count the first year when adding, it was already counted. They want the "total" amount of salaries, meaning to add all of them up to get the overall salary, or money earned for 10 years. Sorry for the long message, might make it more confusing o_O

OpenStudy (anonymous):

Brb

OpenStudy (anonymous):

blah. I don't understand.

OpenStudy (anonymous):

Yeah too complicated like that

OpenStudy (anonymous):

Wonder if I did something wrong...

OpenStudy (jim766):

typically if all they wanted to know was how much did the guy make in the tenth year, they would say the very specifically....how much did he make in year 10 or the 10th year. Instead the chose to say find the "total salary over the ten year period"

OpenStudy (jim766):

that says to me to take all of his yearly salaries and add them together for the first 10 years.

OpenStudy (anonymous):

Thank you, that was what I've been trying to say, just couldn't get it out

OpenStudy (anonymous):

The proper way*

Directrix (directrix):

The 10th term is: $37 000 + (10 - 1) * 2 500 = 59 500. The sum of the first ten terms is: 10*( 37 000 + 59 500 ) / 2 = $ 482 500 --> same as @Jim766 Link to formulas: http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

okay way to go @Jim766 save the day!!!

OpenStudy (anonymous):

Guess we're done here

OpenStudy (anonymous):

yes sir, Thank you all for helping. I appreciate it.

OpenStudy (anonymous):

I tried to, your welcome

OpenStudy (anonymous):

You did help

OpenStudy (anonymous):

Thanks for the compliment :)

OpenStudy (anonymous):

Jim did a good job